OpenStudy (deoxna):

Given two non-zero vectors a and b such that [a+b]=[a-b], find the values of a*b. (Where "[" inidcates magnitudes and "*" is the dot product)

OpenStudy (raden):

looks a*b must be 0

OpenStudy (deoxna):

That's what I was thinking by playing on geogebra, the magnitudes are equal when the two vector are perpendicular. But I'm not sure how to prove or confirm that.

OpenStudy (raden):

[a+b] = sqrt{(a1+b1)^2+(a2+b2)^2+(a3+b3)^2} [a+b] = sqrt{(a1)^2+(a2)^2+(a3)^2+2(a1b1+a2b2+a3b3)+(b1)^2+(b2)^2+(b3)^2} [a+b] = sqrt([a]^2+[b]^2+2a*b) [a-b] = sqrt{(a1-b1)^2+(a2-b2)^2+(a3-b3)^2} [a-b]= sqrt{(a1)^2+(a2)^2+(a3)^2+2(a1b1+a2b2+a3b3)+(b1)^2+(b2)^2+(b3)^2} = sqrt([a]^2+[b]^2-2a*b) because [a+b] = [a-b], so sqrt([a]^2+[b]^2+2a*b) = sqrt([a]^2+[b]^2-2a*b)

OpenStudy (raden):

square to both sides, gives [a]^2+[b]^2+2a*b = [a]^2+[b]^2-2a*b solve for ab

OpenStudy (raden):

4ab = 0 ab = 0/4 = 0

OpenStudy (deoxna):

I don't get what you are doing in the second line of you first comment...

OpenStudy (deoxna):

Forget it, I see you already expanded and simplified the expressions...

OpenStudy (raden):

u just neeed to expand (a1+b1)^2 = ....

OpenStudy (raden):

yea... ^^^

OpenStudy (raden):

also (a1-b1)^2 = ....

OpenStudy (raden):

ops... miss index of a and b

OpenStudy (raden):

(a+b)^2 and (a-b)^2 i meant

OpenStudy (deoxna):

But in the third line, how do you simplify it to [a]^2?

OpenStudy (raden):

just remember that define of [a] = sqrt{(a1)^2+(a2)^2+(a3)^2}, square to both sides gives [a]^2 = (a1)^2+(a2)^2+(a3)^2

OpenStudy (raden):

also [b] = sqrt{(b1)^2+(b2)^2+(b3)^2}, square to both sides gives [b]^2 = (b1)^2+(b2)^2+(b3)^2

OpenStudy (raden):

and a*b = a1b1 + a2b2 + a3b3

OpenStudy (raden):

does this make sense ?

OpenStudy (raden):

just compare this sqrt(4) = 2 square both sides, be 4 = 4

OpenStudy (deoxna):

Yeah thanks! I had done the same in two dimensions, but was missing the jump from a1b1+a2b2=a*b-0. Anyway thanks for all the help.

OpenStudy (deoxna):

Sorry, meant to say: a1b1+a2b2=a*b=0

OpenStudy (raden):

yeah... np very welcome